Please help me solve this problem:
y=1/2x-3
y=1/3x+2
I tried this
y=1/3x+2
-1/3x+y=2
-y=-y
-1/3x =-y+2
3/1(-1/3x)=(-y+2)3/1
x=-3y+6
Then substitue x=-3y+6
y=1/x(-3y+6)-3
y=-3/xy+6/2-3
y=-3/2y+3-3

But now I am lost

Answer by stanbon(57984)

y=1/2x-3
y=1/3x+2
------------
Substitute to get:
(1/2)x - 3 = (1/3)x + 2
[(1/2)-(1/3)]x = 5
(1/6)x = 5
x = 30
-----------
y = (1/2)30 - 3
y = 12
======================
Cheers,
Stan H.